OpenStudy (anonymous):

Simplify 5 square root of 11 end root minus 12 square root of 11 end root minus 2 square root of 11 A. negative 33 square root of 9 B. negative 11 square root of 9 C. negative 9 square root of 11 D. negative 9 square root of 33

OpenStudy (anonymous):

\[5\sqrt{11} - 12\sqrt{11} - 2\sqrt{11}\]

OpenStudy (marihelenh):

Because they all have the same square roots, they are like terms, so you can subtract them all together.

OpenStudy (mathstudent55):

Let's say a positive number of dollars is money you have, and a negative number of dollars is money you owe. What is $5 - 12$ - $2 = ?

OpenStudy (anonymous):

-9

OpenStudy (marihelenh):

Yep, and then the square root stays the same because they are like terms.

OpenStudy (mathstudent55):

Correct. You'd have -$9 It's the same with x and with \(\sqrt{11} \) Since you are combining a number of the same things, in this case all terms have \(\sqrt {11} \), just combine the coefficients and then write \(\sqrt{11} \) at the end. $5 - $12 - $2 = -$9 5x - 12x - 2x = -9x Similarly: \(5\sqrt{11} - 12\sqrt{11} - 2\sqrt{11} = -9\sqrt {11}\)

OpenStudy (anonymous):

Thank you very much sir!

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

Can you help me with 1 more?